OpenStudy (anonymous):

What are intervals of increase and decrease for f(x)=7+6x-x²

OpenStudy (anonymous):

\[\huge f'(x)= 6-2x\] \[\huge 0=6-2x\] \[\huge 3=x\]

OpenStudy (anonymous):